Lösung der Aufgabe 1:

geg: f(x) = (z − 2y, −3z + 2y − 3x, z − y)

Basis: e1 = ( \begin{matrix}1 \\ 0 \\ 0 \end{matrix} )  e2 = ( \begin{matrix}1 \\ 0 \\ 1 \end{matrix} )  e3 = ( \begin{matrix}0 \\ 1 \\ 1 \end{matrix} )

Berechnung der Darstellungsmatrix

f(e1) = ( \begin{matrix}0 \\ −3 \\ 0 \end{matrix} ) = −3 · e1 + 3 · e2 − 3 · e3 f(e2) = ( \begin{matrix}1 \\ 0 \\ 1 \end{matrix} ) = 0 · e1 + 1 · e2 + 0 · e3

f(e3) = ( \begin{matrix}−1 \\ −1 \\ 0 \end{matrix} ) = −2 · e1 + 1 · e2 − 1 · e3

A_{BB} = ( \begin{matrix}−3 & 0 & −2 \\ 3 & 1 & 1 \\ −3 & 0 & −1 \end{matrix} )

Lösung:

Damit ergeben sich als invariante Unterräume: U_{1} = (e_{2}) und U_{2} = (e_{1}, e_{2}, e_{3})